THE JET-COOLED HIGH RESOLUTION Ã^{2}E′′-X̃^{2}A′_{2} VIBRONIC BANDS OF NO_{3}

\underline{MING-WEI CHEN}, TERRANCE J. CODD, GABRIEL M. P. JUST^{a}, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

The 0_{0}^{0}, 4_{0}^{1}, 4_{0}^{2} and 2_{0}^{1} vibronic bands of the Ã state NO_{3} absorption spectrum has been successfully observed with our high-resolution, jet-cooled cavity ring-down apparatus. Ground state combination differences^{b}^{c} are used to analyze all four vibronic bands. Rotational transitions of the 4_{0}^{1} band (parallel band) band are assigned including some levels that appear to be doubled. The 4_{0}^{2} band verifies the existence of anomalous doublets in both the 4_{0}^{1} and 4_{0}^{2} bands. The 2_{0}^{1} band is a perpendicular band which has a different band type. Preliminary assignments of this band are utilized for the comprehensive understanding of the structure of NO_{3} in the Ã state. Besides the ν_{2} and ν_{4} vibronic bands, the vibronically forbidden origin band (0_{0}^{0} band) has also been recorded. The weakly observed Ã-X̃ origin band structure appears to be different from either the parallel or perpendicular band type and is likely a magnetic dipole transition.

apresent address: Lawrence Berkeley National Laboratory, Berkeley, CA 94720

bK. Kawaguchi, E. Hirota, T. Ishiwata, and I. Tanaka, J. Chem. Phys., \underline{93}, 951 (1990)

cK. Kawaguchi, T. Ishiwata, E. Hirota, and I. Tanaka, Chem. Phys., \underline{231}, 193 (1998)